OpenStudy (anonymous):

help The sum or product of a rational number and an irrational number is always rational. irrational. a repeating decimal. a fraction.

OpenStudy (anonymous):

Look at this and answer : \[\sqrt{2}+1/2\] Can you answer this ?:)

OpenStudy (anonymous):

1.9 i converted it

OpenStudy (anonymous):

can you answer to \[\sqrt{2}=?\]

OpenStudy (anonymous):

1.4

OpenStudy (anonymous):

heh ! 1.4*1.4=?

OpenStudy (anonymous):

1.59 n keeps going

OpenStudy (anonymous):

i mean 1.95

hero (hero):

An irrational number is has infinite decimal places. Adding a terminating decimal to it will not make it rational.

OpenStudy (anonymous):

so its because the product is always repeating or terminating decimal?

hero (hero):

\(\sqrt{2} = 1.414213562373095048801688724209698078569671875376948073176679...\) Add .5 to that and you will get \(1.914213562373095048801688724209698078569671875376948073176679...\) As you can see, the sum is still irrational.

OpenStudy (anonymous):

oooooooooooooooooooooooo ok i see thank you.